Stella can type 110 words in 2 1/2 minutes. At the same rate, how many words can she type in 6 minutes?
Answer by stanbon(57337)

Stella can type 110 words in 2 1/2 minutes. At the same rate, how many words can she type in 6 minutes?
---------------------------------
Proportion:
rate = 110/(5/2) = 2*22 = 44 words per minute
----
# of words in 6 minutes = 6*44 = 264 words
-----------------
Cheers,
Stan H.
=================